OpenStudy (filllip66):

Kraken and seven other pirates agreed to share a pile of gold pieces equally. Right before they divided up the gold, two more pirates showed up and demanded to be included, so the gold ended up getting divided among ten pirates instead. Kraken received 40 fewer pieces of gold than he would have if the two extra pirates hadn't arrived. How many pieces of gold did Kraken receive?

OpenStudy (tteesstt8080):

Okay, so what would YOU do first? I would make two equations.

OpenStudy (calculusxy):

@Filllip66 It's true that you may need to make two equations for this problem.

OpenStudy (filllip66):

you do 8a=10b and b=a-40

OpenStudy (calculusxy):

Make one equation where x is the pieces of gold that is to be divided equally and y is the number of pieces that Kraken would receive if there were only 8 pirates. Make the second equation with the same x and y variables but now if there were only 10 pirates.

OpenStudy (filllip66):

So I got the answer a=200

OpenStudy (filllip66):

But I need B

OpenStudy (calculusxy):

What does a and b mean?

OpenStudy (calculusxy):

You need to properly declare your variables.

OpenStudy (filllip66):

Oh I got the answer

OpenStudy (filllip66):

Thank you

OpenStudy (filllip66):

It equals 160

OpenStudy (filllip66):

Thank you

OpenStudy (calculusxy):

Those are the pieces of gold that Kraken if there were 8 and 10 pirates. But what is the original number of pieces to be divided equally?

OpenStudy (filllip66):

160

OpenStudy (calculusxy):

Incorrect.

OpenStudy (filllip66):

If not 200

OpenStudy (filllip66):

I think

OpenStudy (calculusxy):

That is also incorrect.

OpenStudy (calculusxy):

I am pretty sure that you are very close to the actual answer.

OpenStudy (filllip66):

Hmmmm.............

OpenStudy (filllip66):

DO NOT KNOW

OpenStudy (calculusxy):

Here are my equations: \(x\) = pieces of gold to be divided \(y\) = pieces of gold that Kraken will receive with 10 pirates for 8 pirates... \[\frac{x}{8} = 40 + y\] for 10 pirates... \[\frac{x}{10} = y\]

OpenStudy (calculusxy):

Do you understand how I came about writing my equations?

OpenStudy (filllip66):

Yes

OpenStudy (calculusxy):

Now use elimination or substitution (although I prefer the latter method) to solve for the value of \(x\).

OpenStudy (filllip66):

Dont get it

OpenStudy (calculusxy):

Are you stuck with solving for x?

OpenStudy (filllip66):

I WAS RIGHT

OpenStudy (calculusxy):

What do you mean?

OpenStudy (filllip66):

I CHECKED AND GOT IT

OpenStudy (filllip66):

But still thank you

OpenStudy (calculusxy):

Shoot! I did not understand the directions very clearly. Very sorry :(

OpenStudy (filllip66):

Hey